OpenStudy (anonymous):

Is this factorable? x is a natural number a,b are real numbers a^2 + b^2 a^4 + b^4 and more generally: a^(2x) + b^(2x) Because, every polynomials should have real roots, right?

OpenStudy (mayankdevnani):

@pourushgupta

hartnn (hartnn):

nopes, polynomials can have imaginary roots. a^2+b^2 can be factored as (a+ib)(a-ib)

hartnn (hartnn):

a^4+b^4 can be factored as \(a^4+2a^2b^2+b^4-2a^2b^2 = (a^2+b^2)^2-(\sqrt 2ab)^2\) then using x^2-y^2 formula

OpenStudy (anonymous):

what about the general case?

hartnn (hartnn):

i cannot think of any way to factor that general case, but u can find the root(if u consider that as equation in x), by equating it to 0, it will be imaginary root.

OpenStudy (unklerhaukus):

\[(a^2 + b^2)=(a+ib)(a-ib)\] \[(a^4 + b^4)=(A^2+B^2)\]\[=(A+iB)(A-iB)=(a^2+ib^2)(a^2-ib^2)\]